# DEGAP: Dual Event-Guided Adaptive Prefixes for Templated-Based Event Argument Extraction Model with Slot Querying

### v1

@@ Abstract
Recent advancements in event argument extraction (EAE) involve incorporating beneficial auxiliary information into models during training and inference, such as retrieved instances and event templates. Additionally, some studies introduce learnable prefix vectors to models. These methods face three challenges: (1) insufficient utilization of relevant event instances due to deficiencies in retrieval; (2) neglect of important information provided by relevant event templates; (3) the advantages of prefixes are constrained due to their inability to meet the specific informational needs of EAE. In this work, we propose DEGAP, which addresses the above challenges through two simple yet effective components: (1) dual prefixes, where the instance-oriented prefix and template-oriented prefix are trained to learn information from different event instances and templates, respectively, and then provide relevant information as cues to EAE model without retrieval; (2) event-guided adaptive gating mechanism, which guides the prefixes based on the target event to fully leverage their advantages. Extensive experiments demonstrate that our method achieves new state-of-the-art performance on four datasets (ACE05, RAMS, WIKIEVENTS, and MLEE). Further analysis verifies the importance of the proposed design and the effectiveness of the main components [^1] We will release the code after acceptance..
@@ 1 Introduction
As an essential subtask of Event Extraction, Event Argument Extraction (EAE) aims to identify event arguments and their roles for given event mentions in context. For example, in Fig. 1, given a Life.Die event with trigger killed, EAE models need to extract arguments soldiers, bomber, and checkpoint, and the corresponding roles — Victim, Agent, and Place. In the field of NLP, EAE has always been a challenging task, which requires deep natural language understanding to address Hsu et al. (2023).
Consequently, some efforts incorporate ancillary information beneficial for EAE into models during training and inference, such as augmenting input with retrieved instances Du and Ji (2022); Ren et al. (2023); Zhao et al. (2023a) or querying arguments using role slots in event template Liu et al. (2024); Zhang et al. (2022); Ma et al. (2022); He et al. (2023); Nguyen et al. (2023). Moreover, in light of advancements in prompt learning Li and Liang (2021); Liu et al. (2022b), some studies introduce prefixes (i.e, learnable prompt vectors) into models Liu et al. (2022a); Zhang et al. (2023a); Hsu et al. (2023); Nguyen et al. (2023). Despite the empirical success of previous works, three issues remain that have not been appropriately solved, hindering further performance improvements.
@@
First, relevant event instances are not fully and accurately utilized due to deficiencies in retrieval. Previous studies regard relevant instances retrieved from training data Wang et al. (2022) as cues to extract arguments and learn by analogy. However, due to the need to pre-specify the number of event instances to be retrieved and potentially irrelevant instances retrieved Ren et al. (2023), these methods struggle to fully and accurately utilize relevant reference instances to boost performance. Additionally, these methods often suffer from high computational costs and nontrivial knowledge loss because they independently retrieve and encode event instances, resulting in a poor local optimum Huang et al. (2023b). Here raises an interesting question: Can EAE models benefit from relevant event instances without relying on retrieval?
@@
Second, relevant event templates are not considered. Merely using the corresponding template for the target event as prompt, some important information provided by relevant event templates might be overlooked Ma et al. (2022); He et al. (2023). Moreover, only Ren et al. (2023) attempted to retrieve role labels relevant to the target event as demonstrations and achieved success. As shown in Fig. 1, event templates can provide rich semantic information about events, and relevant events often have similar structures and share partial roles (e.g., Life.Die and Conflict.Attack) Li et al. (2023a). Intuitively, regarding relevant event templates as cues and learning by analogy can help EAE models better extract the target event from complex semantics.
Finally, existing prefixes largely fail to meet the specific informational needs of EAE, leading to unsatisfactory performance. Typically, existing prefixes used for EAE are generalizable across multiple tasks, rather than specifically tailored to a particular task. Despite high portability, recent research indicates that these prefixes do not effectively meet the specific informational needs of the task, failing to fully utilize the advantages of prefix Zhang et al. (2023d). Meanwhile, for the EAE task, some studies show that models should focus on information relevant to the target event and irrelevant information would hinder the model’s performance Xu et al. (2022); Liu et al. (2023c). However, how to adaptively capture relevant information from the prefix at each layer based on the target event, and how to guide the prefix to leverage its advantages fully, remain largely under-explored.
To address these challenges, we propose DEGAP (Dual Event-Guided Adaptive Prefixes for event argument extraction). Inspired by recent studies that utilize prefixes as efficient parameters to learn information from different corpora and integrate retrieved instances into prefixes as cues Zhang et al. (2023b); Liu et al. (2023a), we design two types of prefixes: The instance-oriented prefix is trained to learn event instance information at the event mention-level, and the template-oriented prefix is trained to learn event template information at the ontology level. To adaptively capture relevant information from the prefix at each layer based on the target event and further guide the prefix, we design an event-guided adaptive gating mechanism. It utilizes the target event’s trigger and type to score each instance-oriented and template-oriented prefix token via gated weight assignment at each layer, respectively. Finally, the event-guided prefixes provide relevant information as cues at each layer of the model during inference without retrieval, facilitating the model to extract the target event.
@@ 2 Related Work
Event argument extraction. Traditionally, most models are classification-based Chen et al. (2015); Nguyen et al. (2016); Yang et al. (2018); Ebner et al. (2020); Zhang et al. (2020); Du and Cardie (2020a); Xu et al. (2022). Recently, there has been a trend towards framing EAE as a text generation or question-answering (QA) task. Generation-based methods sequentially generate all arguments of the target event by following a pre-defined template Li et al. (2021); Hsu et al. (2021); Zeng et al. (2022); Hsu et al. (2023). QA-based methods query the arguments of each role using corresponding question templates Du and Cardie (2020b); Liu et al. (2021); Wei et al. (2021); Lu et al. (2023); Uddin et al. (2024) or role slots in event template Zhang et al. (2022); Ma et al. (2022); Li et al. (2023a); He et al. (2023).
This work builds on the existing state-of-the-art (SOTA) methods PAIE Ma et al. (2022) and TabEAE He et al. (2023), which solves EAE by using role slots in the event template to query input context for extracting argument spans. With two proposed simple yet effective core components—the dual prefixes and the event-guided adaptive gating mechanism—our method establishes new SOTA performance on four datasets.
Event extraction with auxiliary information. As a task requiring a deep understanding of context, many previous efforts have been put into exploring which auxiliary information is useful for event predictions. One line of studies is incorporating such auxiliary information during training and inference, such as event templates Li et al. (2021); Hsu et al. (2021); Zeng et al. (2022); He et al. (2023); Liu et al. (2024), natural questions Liu et al. (2020, 2021); Lu et al. (2023); Uddin et al. (2024), retrieved instances Du and Ji (2022); Ren et al. (2023); Zhao et al. (2023a), already predicted events Huang et al. (2023a); Du et al. (2022), and syntactic and semantic information Huang et al. (2017); Veyseh et al. (2020); Ahmad et al. (2021); Xu et al. (2022); Yang et al. (2023). However, as mentioned in Section 1, existing methods have not fully exploited the benefits of relevant event instances and event templates.
Prompt learning. Prompt learning is a straightforward and effective method that adjusts PLMs for specific downstream tasks by adding prompts to the input. One line of research applies an automated search to find appropriate discrete prompting words Wallace et al. (2019); Shin et al. (2020); Wen et al. (2024). Another line of research attaches a set of trainable prefix tokens to the model as soft prompts Li and Liang (2021); Liu et al. (2022b); Chen et al. (2023); Li et al. (2023b), which have been widely applied in various studies. Liu et al. (2023a) integrates retrieved instances into the prefix as cues for inference. Zhang et al. (2023b) uses prefixes to learn information from different corpora and refine inference. Additionally, some EAE efforts use prefixes to optimize manually designed prompts Liu et al. (2022a); Nguyen et al. (2023) or to integrate syntactic and semantic information Hsu et al. (2023). Despite the successes, the prefixes used are not specifically tailored for the EAE task, hence the benefits of prefixes can still be further exploited Zhang et al. (2023d).
@@ 3.1 Model Architecture
Figure 2: The overview of DEGAP. Here, the two encoders share the same parameters, and the two decoders also share the same parameters. Due to space constraints, the event template here is simplified. When encoding the context, $\text{Prefix}_{i}$ and $h_{i-1}$ refer to $\text{Prefix}_{i}^{\textit{ins}}$ and $h_{i-1}^{trig}$, respectively. Conversely, they refer to $\text{Prefix}_{i}^{\textit{tem}}$ and $h_{i-1}^{type}$, respectively. IOP: Instance-Oriented Prefix. TOP: Template-Oriented Prefix.
Fig. 2 shows the overall architecture of DEGAP. For given event type $e$, we create an event template for it (§ 3.1.1). To learn and save information from event instances and event templates, we design instance-oriented prefixes and template-oriented prefixes, respectively (§ 3.1.2). Then, we design an event-guided adaptive gating mechanism to adaptively capture relevant information in the prefixes and further guide the prefixes (§ 3.1.3). Subsequently, when encoding the context and event template, these event-guided prefixes provide relevant event information as cues to the model at each layer without retrieval (§ 3.1.4). Finally, we introduce the span selection, which utilizes role slots in the event template to query the context for extracting argument spans (§ 3.1.5).
@@ 3.1.1 Event Template Creation
We create an event template $\mathcal{T}_{e}$ for the given event type $e$, utilizing its rich semantic information to guide the model Hsu et al. (2022). Similar to previous works Liu et al. (2022a); Zhang et al. (2023b), the event template $\mathcal{T}_{e}$ consists of two parts: the type part $\mathcal{I}_{e}$ and the schema part $\mathcal{S}_{e}$. The type part $\mathcal{I}_{e}$ adopts the paradigm “Event type is [Event type]”. For the schema part $\mathcal{S}_{e}$, we directly inherit it from TabEAE He et al. (2023). Considering the example in Fig. 2, where there is a Conflict.Attack event with the event trigger hit, the template for this event is: “<s> Event type is Conflict.Attack </s> <s> Attacker (and Attacker) attacked Target (and Target) using Instrument (and Instrument) at Place (and Place) </s>”, where each underlined part (i.e., role slot) is named after an argument role. The brackets indicate that there may be multiple arguments of that role in an event.
@@ 3.1.2 Dual Prefixes
To fully benefit from relevant event instances and event templates without retrieval, we construct two types of prefixes: (1) the instance-oriented prefix and (2) the template-oriented prefix. With self-attention mechanisms, these prefixes can learn rich semantic information about events from different instances and templates and further guide the model to extract arguments of the target event.
Instance-Oriented Prefix (IOP). To learn and save information from different event instances, at the $i$-th layer of the encoder or decoder, we equip the self-attention block with a set of learnable prefix tokens oriented towards event instances $\text{Prefix}_{i}^{\textit{ins}}\in\mathbb{R}^{len_{\textit{ins}}\times m}$ , where $len_{\textit{ins}}$ is the length of $\text{Prefix}_{i}^{\textit{ins}}$ and $m$ corresponds to the hidden dimension.
Template-Oriented Prefix (TOP). To learn and save information from different event templates, at the $i$-th layer of the encoder or decoder, we equip the self-attention block [^2] We also attempt to equip the cross-attention block in each layer of the decoder with the prompt-oriented prefix, but observe almost no improvement in performance. with a set of learnable prefix tokens oriented towards event templates $\text{Prefix}_{i}^{\textit{tem}}\in\mathbb{R}^{len_{\textit{tem}}\times m}$, where $len_{\textit{tem}}$ is the length of $\text{Prefix}_{i}^{\textit{tem}}$.
@@ 3.1.3 Event-Guided Adaptive Gating mechanism
Not all information within the prefix benefits argument extraction of the target event. To adaptively capture relevant information from the prefix at each layer based on the target event and further guide the prefix, inspired by Liu et al. (2023b), we utilize the target event’s trigger and type to score each prefix token in $\text{Prefix}_{i}^{ins}$ and $\text{Prefix}_{i}^{tem}$ through gated weight assignment, respectively. For example, in the $i$-th layer, the weight distribution of tokens in $\text{Prefix}_{i}^{ins}$ is calculated as:
@@
where $h_{i-1}^{trig}$ is the hidden state of the event trigger from the previous layer obtained through average pooling, and $W^{ins}_{i}\in\mathbb{R}^{m\times len_{ins}}$ corresponds to the parameters to be learned.
Additionally, we also follow Zhang et al. (2023d) to introduce a learnable scaled factor $\lambda_{i}^{ins}$ to balance the information provided by the original input and the prefix. We obtain the final instance-oriented prefix representation $\text{Prefix}_{i}^{ins^{\prime}}$:
@@
where $\odot$ is the element-wise multiplication. Similarly, we derive the finally template-oriented prefix representation $\text{Prefix}_{i}^{tem^{\prime}}$:
@@
where $h_{i-1}^{type}$ is the hidden state obtained through average pooling of the event type from the previous layer, and $W^{tem}_{i}\in\mathbb{R}^{m\times len_{tem}}$ is the learnable parameters.
Finally, the new self-attention module with event information intervention is formalized as:
@@
where $\text{MHSA}(\cdot)$ denotes the multi-head self-attention mechanism, $\text{Prefix}i^{{}^{\prime}}$ represents either $\text{Prefix}{i}^{ins^{\prime}}$ or $\text{Prefix}{i}^{tem^{\prime}}$, $H$ refers to the hidden state, and $(\cdot){b:c}$ denotes the slicing operation on the $seq_{l}en$ dimension from $b$ to $c$.

### v2

# DEGAP: Dual Event-Guided Adaptive Prefixes for Templated-Based Event Argument Extraction with Slot Querying

@@ Abstract
Recent advancements in event argument extraction (EAE) involve incorporating useful auxiliary information into models during training and inference, such as retrieved instances and event templates. These methods face two challenges: (1) the retrieval results may be irrelevant and (2) templates are developed independently for each event without considering their possible relationship. In this work, we propose DEGAP to address these challenges through a simple yet effective components: dual prefixes, i.e. learnable prompt vectors, where the instance-oriented prefix and template-oriented prefix are trained to learn information from different event instances and templates. Additionally, we propose an event-guided adaptive gating mechanism, which can adaptively leverage possible connections between different events and thus capture relevant information from the prefix. Finally, these event-guided prefixes provide relevant information as cues to EAE model without retrieval. Extensive experiments demonstrate that our method achieves new state-of-the-art performance on four datasets (ACE05, RAMS, WIKIEVENTS, and MLEE). Further analysis shows the impact of different components.
@@ 1 Introduction
As an essential subtask of Event Extraction, Event Argument Extraction (EAE) aims to identify event arguments and their roles for given event mentions in context. For example, in Fig. 1, given a Life.Die event with trigger killed, EAE models need to extract arguments soldiers, bomber, and checkpoint, and the corresponding roles — Victim, Agent, and Place. In the field of NLP, EAE has always been a challenging task, which requires deep natural language understanding Hsu et al. (2023).
Consequently, some efforts incorporate ancillary information useful for EAE into models during training and inference, such as augmenting input with retrieved instances Du and Ji (2022); Ren et al. (2023); Zhao et al. (2023a) or querying arguments using role slots in event template Liu et al. (2024); Ma et al. (2022); He et al. (2023); Nguyen et al. (2023). Moreover, in light of advancements in prompt learning Li and Liang (2021); Liu et al. (2022b), some studies introduce prefixes (i.e, learnable prompt vectors) into models as additional keys and values to influence the computation of attention Liu et al. (2022a); Zhang et al. (2023a); Hsu et al. (2023); Nguyen et al. (2023). Despite the empirical success of previous studies, two issues remain, hindering further performance improvements.
@@
First, retrieved instances are not always reliable due to deficiencies in retrieval. Previous studies regard event instances retrieved from training data Wang et al. (2022) as cues to extract arguments. However, these methods typically use a retriever pre-trained on other tasks with frozen parameters, which may not always retrieve other event instances relevant to the target event  [^1] In a context, only a few words are event arguments, while other distracting context can mislead retrieval based on sentence semantic similarity. According to Ren et al. (2023), on the RAMS dataset Ebner et al. (2020), only 16.51% of event instances can recall a sample with the same event type. . Additionally, these methods often suffer from high computational costs and nontrivial knowledge loss because they independently retrieve and encode event instances, resulting in a poor local optimum. Directly increasing the number of instances to be retrieved is typically inefficient. Recent advancements show that one can utilize prefixes as efficient parameters to learn information from different corpora and integrate retrieved instances into prefixes as cues for inference Zhang et al. (2023b); Liu et al. (2023a). Inspired by this intriguing observation, in this work, we attempt to explore a novel approach to utilize relevant event instance without retrieval. First, we use prefixes to learn the semantic information of event instances. Then, we leverage the learned prefixes as cues to guide the model.
@@
Second, no relation between different event templates is considered. Most of the existing approaches use a separate template as prompt for each target event. However, some event templates for other relevant events may provide useful information to the given event. This potential connection between event templates has been overlooked in the literature Ma et al. (2022); He et al. (2023). Only Ren et al. (2023) attempted to retrieve role labels from other events that are relevant to the target event as demonstrations. As shown in Fig. 1, event templates provide rich semantic information about events, and other relevant events often have similar structures and share part of their roles (e.g., Life.Die and Conflict.Attack) Li et al. (2023). Intuitively, one can regard relevant event templates as cues and perform reasoning by analogy. This can help EAE models better extract the target event from complex semantics.
In this paper, we aim to address the above challenges. We propose DEGAP (Dual Event-Guided Adaptive Prefixes for event argument extraction). Specifically, we design two types of prefixes: the instance-oriented prefix and the template-oriented prefix. The reason we design separate prefixes for event instances and templates, rather than a shared prefix for both, is that instances and templates contain different types of information about events: an instance is a concrete example of an event, while a template tries to capture the general structure of an event. They play different roles in EAE, thus it is more appropriate to train separate prefixes for them [^2] Our experimental results also confirm that dual prefixes can better learn semantic information from different event instances and templates, and further guide model to extract argument (§ 5.2).. These instance/template-specific parameters are used to learn information from different event instances/templates and enable them to interact sufficiently with each other, rather than ignoring the potential connections between them. They are trained and updated together with model parameters, thus avoiding the insufficient retrieval and poor local optimum issues present in retrieval-based EAE methods. Moreover, not all event instances and templates are useful to the target event. We aim to use only those instances and templates that are relevant to the target event Xu et al. (2022); Liu et al. (2023c). To connect other relevant events to the target event and thus adaptively capture relevant information from the prefix, we design an event-guided adaptive gating mechanism. It utilizes the target event’s trigger and type to score each instance-oriented and template-oriented prefix token via gated weight assignment at each layer, respectively. Finally, these event-guided prefixes are used to guide model for extracting argument.
@@ 2 Related Work
Event (argument) extraction with auxiliary information. As the task requires a deep understanding of context, many previous efforts have been put into exploring which auxiliary information is useful for event predictions. One line of studies is incorporating such auxiliary information during training and inference, such as event templates Li et al. (2021); Hsu et al. (2021); Zeng et al. (2022); He et al. (2023); Liu et al. (2024), natural questions Liu et al. (2020, 2021); Lu et al. (2023); Uddin et al. (2024), retrieved instances Du and Ji (2022); Ren et al. (2023); Zhao et al. (2023a), already predicted events Huang et al. (2023); Du et al. (2022), and syntactic and semantic information Huang et al. (2017); Veyseh et al. (2020); Ahmad et al. (2021); Xu et al. (2022); Yang et al. (2023). As mentioned in § 1, these methods may lead to irrelevant results due to retrieval deficiencies that may negatively impact the final results. Our approach is based on the principle of prompt learning.
Prompt learning. Prompt learning is a straightforward and effective method that adjusts PLMs for specific downstream tasks by adding prompts to the input. One line of research applies an automated search to find appropriate discrete prompting words Wallace et al. (2019); Shin et al. (2020); Wen et al. (2024). An alternative line of research attaches a set of trainable prefix tokens to the model’s attention module as additional keys and values Li and Liang (2021); Liu et al. (2022b); Chen et al. (2023). This method has been widely applied in various studies. For example, Liu et al. (2023a) integrates retrieved instances into the prefix as cues for inference. Zhang et al. (2023b) uses prefixes to learn information from different corpora and refine inference. Additionally, some EAE efforts use prefixes to optimize manually designed prompts Liu et al. (2022a); Nguyen et al. (2023) or to integrate syntactic and semantic information Hsu et al. (2023).
Inspired by these advancements Zhang et al. (2023b); Liu et al. (2023a), We explore a novel approach to utilize potentially relevant event information rather than retrieval: using prefix to learn the semantic information of different event instances and templates, then leveraging the learned prefix as cues to guide model. Moreover, previous methods overlook the potential connections between event templates. Some relevant event templates may provide useful information to the target event. As such, we propose an event-guided adaptive gating mechanism, which can effectively connect other relevant events to the target event and thus adaptively capture relevant information from the prefix.
@@ 3.1 Model Architecture
Figure 2: The overview of DEGAP. Here, the two encoders share the same parameters, and the two decoders also share the same parameters. Due to space constraints, the event template here is simplified. When the model input is context, $\text{Prefix}_{i}$ and $h_{i-1}$ refer to $\text{Prefix}_{i}^{\textit{ins}}$ and $h_{i-1}^{trig}$, respectively. Conversely, they refer to $\text{Prefix}_{i}^{\textit{tem}}$ and $h_{i-1}^{type}$, respectively. IOP: Instance-Oriented Prefix. TOP: Template-Oriented Prefix.
Fig. 2 shows the overall architecture of DEGAP. For a given event type $e$, we create an event template for it (§ 3.1.1). To learn information from event instances and templates, we design instance-oriented prefixes and template-oriented prefixes, respectively (§ 3.1.2). Dual prefixes can better learn semantic information from different event instances and templates compared to a shared prefix (§ 5.2). Then, we design an event-guided adaptive gating mechanism to connect other relevant events to the target event and thus adaptively capture relevant information in the prefix (§ 3.1.3). Subsequently, these event-guided prefixes provide relevant event information as cues to model (§ 3.1.4). Finally, we introduce the span selection, which utilizes role slots in the event template to query context for extracting argument (§ 3.1.5).
@@ 3.1.1 Event Template Creation
We create an event template $\mathcal{T}_{e}$ for the given event type $e$, utilizing its rich semantic information to guide the model Hsu et al. (2022). Similar to previous works Liu et al. (2022a); Zhang et al. (2023b), the event template $\mathcal{T}_{e}$ consists of two parts: the type part $\mathcal{I}_{e}$ and the schema part $\mathcal{S}_{e}$. The type part $\mathcal{I}_{e}$ adopts the format “Event type is [Event type]”. For the schema part $\mathcal{S}_{e}$, we directly use that from TabEAE He et al. (2023). Considering the example in Fig. 2, where there is a Conflict.Attack event with the event trigger hit, the template for this event is: “<s> Event type is Conflict.Attack </s> <s> Attacker (and Attacker) attacked Target (and Target) using Instrument (and Instrument) at Place (and Place) </s>”, where each underlined part (i.e., role slot) is named after an argument role. The brackets indicate that there may be multiple arguments of that role in an event.
@@ 3.1.2 Dual Prefixes
To directly and fully benefit from relevant event instances and templates without retrieval as an intermediate step, we construct two types of prefixes: (1) the Instance-Oriented Prefix (IOP) and (2) the Template-Oriented Prefix (TOP). These instance/template-specific prefix parameters can learn rich semantic information from different event instances/templates and are trained and updated alongside model parameters.
Note that we only design a shared IOP/TOP for all event instances/templates, rather than one for each event instance/template. This approach allows them to interact fully with each other, rather than ignoring the potential connections between them.
To be more specific, at the $i$-th layer of encoder or decoder, we equip the self-attention block with a set of learnable prefix tokens oriented towards event instances, $\text{Prefix}_{i}^{\textit{ins}}\in\mathbb{R}^{len_{\textit{ins}}\times m}$, and event templates, $\text{Prefix}_{i}^{\textit{tem}}\in\mathbb{R}^{len_{\textit{tem}}\times m}$. Here, $len_{\textit{ins}}$ is the length of $\text{Prefix}_{i}^{\textit{ins}}$, $len_{\textit{tem}}$ is the length of $\text{Prefix}_{i}^{\textit{tem}}$, and $m$ corresponds to the hidden dimension. We also attempt to equip the cross-attention block in each layer of the decoder with these prefix tokens, but observe almost no improvement in performance. These additional key and value matrices will be concatenated with the original key and value matrices in the attention block. When calculating dot-product attention, the query at each position will be influenced by these prefix tokens.
@@ 3.1.3 Event-Guided Adaptive Gating mechanism
Some relevant events may provide useful cues to the target event and irrelevant event information might hinder model performance. To connect other relevant events to the target event and thus adaptively capture relevant information from the prefix, inspired by Liu et al. (2023b), we utilize the target event’s trigger and type to score each prefix token in $\text{Prefix}_{i}^{ins}$ and $\text{Prefix}_{i}^{tem}$ through gated weight assignment, respectively. For example, in the $i$-th layer, the weight distribution of tokens in $\text{Prefix}_{i}^{ins}$ is calculated as:
@@
where $h_{i-1}^{trig}$ is the hidden state of the event trigger from the previous layer obtained through average pooling, and $W^{ins}_{i}\in\mathbb{R}^{m\times len_{ins}}$ corresponds to the parameters to be learned. Here, the prefix weight $a_{i}^{ins}$ captures the relevance of other event instances to the given instance, and based on this, we can further capture the relevant information within prefix.
Additionally, we also follow Zhang et al. (2023d) to introduce a learnable scaling factor $\lambda_{i}^{ins}$ to balance the information provided by the original input and the prefix. We obtain the guided instance-oriented prefix representation $\text{Prefix}_{i}^{ins^{\prime}}$:
@@
where $\odot$ is the element-wise multiplication. Similarly, we derive the guided template-oriented prefix representation $\text{Prefix}_{i}^{tem^{\prime}}$:
@@
where $h_{i-1}^{type}$ is the hidden state obtained through average pooling of the event type from the previous layer, and $W^{tem}_{i}\in\mathbb{R}^{m\times len_{tem}}$ is learnable parameters. The prefix weight $a_{i}^{tem}$ captures the relevance of other event templates to the given template.
Finally, these event-guided prefixes can more effectively provide relevant event instances and templates as cues for the target event. The new self-attention module with event information intervention is formalized as:
@@
where $\text{MHSA}(\cdot)$ denotes the multi-head self-attention mechanism, $\text{Prefix}i^{{}^{\prime}}$ represents either $\text{Prefix}{i}^{ins^{\prime}}$ or $\text{Prefix}{i}^{tem^{\prime}}$, $H$ refers to the hidden state, and ${b:c}$ denotes the slicing operation on the $seq\_len$ dimension from $b$ to $c$.